OpenStudy (anonymous):

Please Help Somebody Emergency .. Medal Fan ! A system of equations is shown below: -3x + 7y = -16 -9x + 5y = 16 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)

OpenStudy (anonymous):

i got your back

OpenStudy (anonymous):

-3x + 7y = -16 ---- (1) -9x + 5y = 16 ---- (2) "Replace one equation with the sum of that equation and a multiple of the other." Let us replace (2) by adding (2) to two times (1): 2 times (1): -6x + 14y = -32 Add (2): -9x + 5y = 16 -------------------- -15x + 19y = -16 So the second pair of equations are: -3x + 7y = -16 ---- (1) -15x + 19y = -16 ----- (3) For B) Solve (1) and (2) for x,y Solve (1) and (3) for x,y Show they are the same solutions.

OpenStudy (anonymous):

Thank you so much omg your the best @awesomeness812

OpenStudy (anonymous):

Can you help me with a big favor please I been stuck on this for about a month ?/ @awesomeness812

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

YES